# Feature-aware Hypergraph Generation via Next-Scale Prediction

### v1

@@ 4 Experiments and Results
In this section, we detail our experimental setup, covering datasets, evaluation metrics, results, and ablation studies. The datasets are divided into featureless and featured hypergraphs. For the featureless hypergraphs, we compare FAHNES against HyperPA [26], a Variational Autoencoder (VAE) [27], a Generative Adversarial Network (GAN) [28], and a standard 2D diffusion model [23] trained on incidence matrix images, where hyperedge membership is represented by white pixels and absence by black pixels. For the featured hypergraphs, we compare FAHNES against the graph-based models DiGress [7], DisCo [9], Cometh [29], and DeFoG [30] in a molecular dataset. For 3D meshes, we also compare FAHNES with a sequential disjoint generation baseline (see Figure 1), where we first generate the hypergraph topology and then the features. Extended numerical results and several visualizations of our generated hypergraphs are available in the supplementary material.
The graph-based approaches DiGress, DisCo, Cometh, and DeFoG are only tested on a simple graph dataset because they cannot generate hypergraphs. Although one might attempt to generate hypergraphs using these existing graph generators over the star or clique expansions, this strategy is impractical for several reasons. First, generating the clique expansion of a hypergraph and then recovering the original hyperedges is an NP-hard problem, as it requires enumerating all cliques. Second, in the bipartite representation of a hypergraph, it is non-trivial to determine which partition represents nodes and which represents hyperedges—a distinction that is critical for correct reconstruction. Empirically, we also find that graph-based generative models often fail to produce valid bipartite structures. For instance, we observed that only 30% of outputs from both models [11, 7] generated valid bipartite graphs.
Datasets. We evaluate our method on five featureless datasets: Stochastic Block Model (SBM) [31], Ego [32], Tree [33], ModelNet40 bookshelf, and ModelNet40 piano [34]. We also evaluate FAHNES on three featured hypergraphs datasets, including two sets of 3D meshes and one molecular dataset: Manifold40 bench, Manifold40 airplane [^1] The full set of numerical results for the Manifold40 datasets is provided in the supplementary material. [35], and QM9 [36, 37] with implicit hydrogens. In the case of bench and airplane, node features are 3D positions, and hyperedges do not have features. For QM9, both nodes and hyperedges have features, which are one-hot vectors coding for atom types and bond types, respectively. Although molecules can be represented as traditional graphs, modeling them with hypergraphs opens the door to representing higher-order interactions between atoms. However, in this work, we treat molecules strictly as graphs, with hyperedges corresponding to simple pairwise bonds. Future work could group functional groups’ atoms together in hyperedges. Further details of the datasets are provided in the supplementary material.
Metrics. For featureless hypergraphs, we follow the evaluation criteria in [6]. These include: i) structural comparison metrics such as Node Num (difference in node counts), Node Deg (Wasserstein distance between node degree distributions), and Edge Size (Wasserstein distance between hyperedge size distributions); ii) topological analysis with Spectral (maximum mean discrepancy between the spectral distributions). In scenarios where datasets enforce structural constraints, we report Valid—the proportion of generated samples satisfying those constraints. For all metrics except Valid, lower values indicate improved performance, while higher values are preferable for Valid.
In the case of 3D meshes, we use the nearest training sample Chamfer distance, which computes the Chamfer distance between point clouds sampled from a generated sample and all training samples and outputs the minimum distance. For QM9, we use the same metrics as in [30]: i) overall validity of molecules with Valid (fraction of molecules that can be sanitized using RDKit [38]); ii) diversity of samples with Unique (fraction of unique SMILES generated); and iii) chemical similarity with the learned distribution with FCD (Fréchet inception distance between the activations of ChemNet on the train samples and generated samples [39]).
Implementation details. We adopt the Local PPGN architecture [11, 6]. During generation, we use inpainting to enforce constraints: i) budget splits are fixed to 1 for unexpanded or terminal clusters, ii) equal splits are enforced for expanded clusters with a budget of 2, iii) clusters of size one cannot expand, and iv) features for non-expanded clusters are copied.
@@
Model SBM Hypergraphs Ego Hypergraphs Tree Hypergraphs
Valid SBM $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Valid Ego $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Valid Tree $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$
HyperPA [26] 2.5% 0.075 4.062 0.407 0.273 0% 35.83 2.590 0.423 0.237 0% 2.350 0.315 0.284 0.159
VAE [27] 0% 0.375 1.280 1.059 0.024 0% 47.58 0.803 1.458 0.133 0% 9.700 0.072 0.480 0.124
GAN [28] 0% 1.200 2.106 1.203 0.059 0% 60.35 0.917 1.665 0.230 0% 6.000 0.151 0.459 0.089
Diffusion [23] 0% 0.150 1.717 1.390 0.031 0% 4.475 3.984 2.985 0.190 0% 2.225 1.718 1.922 0.127
HYGENE [6] 65% 0.525 0.321 0.002 0.010 90% 12.55 0.063 0.220 0.004 77.5% 0.000 0.059 0.108 0.012
@@
FAHNES 81.4% 0.010 0.603 0.005 0.005 100% 0.162 0.171 0.129 0.007 82.8% 0.000 0.043 0.046 0.002
Table 2: Evaluation on the ModelNet40 hypergraphs.
Model ModelNet40 Bookshelf ModelNet40 Piano
Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$
HyperPA [26] 8.025 7.562 0.044 0.048 0.825 9.254 0.023 0.067
VAE [27] 47.45 6.190 1.520 0.190 75.35 8.060 1.686 0.396
GAN [28] 0.000 397.2 46.30 0.456 0.000 409.0 86.38 0.697
Diffusion [23] 0.000 20.36 2.346 0.079 0.050 20.90 4.192 0.113
HYGENE [6] 69.73 1.050 0.034 0.068 42.52 6.290 0.027 0.117
@@
FAHNES 0.875 5.600 0.087 0.019 0.125 1.878 0.028 0.036
Table 3: Evaluation on QM9.
Model QM9
Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$
DiGress [7] 99.0% 96.2
DisCo [9] 99.3%
Cometh [29] 99.6% 96.8 0.25
DeFoG [30] 99.3% 96.3 0.12
@@
FAHNES 77.8% 94.3 3.86
Comparison with the baselines. Tables 1 and 3 show the comparisons in the featureless hypergraphs. We observe that FAHNES significantly improves previous methods in many metrics. This is especially evident in the SBM and Ego hypergraphs, where the percentage of valid hypergraphs increases by 16.4% and 10%, respectively. Similarly, the spectral similarity has been improved in the ModelNet40 and Tree hypergraphs.
Budg. Coup. SBM Hypergraphs Ego Hypergraphs Tree Hypergraphs
Valid SBM $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Valid Ego $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Valid Tree $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$
77.7% 0.020 0.889 0.041 0.006 99.0% 2.383 0.239 0.139 0.006 92.2% 0.005 0.066 0.161 0.011
79.9% 0.123 0.815 0.023 0.006 98.5% 1.721 0.223 0.184 0.010 83.8% 0.000 0.045 0.083 0.006
71.6% 0.039 1.413 0.005 0.013 100% 0.459 0.129 0.279 0.010 74.5% 0.005 0.039 0.050 0.003
81.4% 0.010 0.603 0.005 0.005 100% 0.162 0.171 0.129 0.007 82.8% 0.000 0.043 0.046 0.002
Regarding the Manifold40 dataset, FAHNES obtains a Chamfer distance of 0.073 and 0.049 for bench and airplane, respectively, while the sequential baseline reaches 0.143 and 0.117, respectively. This shows the better modeling capacity of FAHNES to jointly generate topology and features in hypergraphs instead of a simple two-step approach. Similarly, Table 3 shows the results in the QM9 dataset, where we observe FAHNES presents promising results, even though our model is not explicitly tailored for the problem of molecule generation. Extended results on featured hypergraphs and visual comparisons of the different approaches are provided in the supplementary material.
@@
Table 5: Ablation studies on the node budget (Budg.) and minibatch OT-coupling (Coup.) for ModelNet40 datasets.
Budg. Coup. ModelNet40 Bookshelf ModelNet40 Piano
Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$
8.6 5.951 0.043 0.027 8.125 3.546 0.075 0.026
7.225 4.183 0.109 0.037 2.375 5.529 0.160 0.028
1.525 5.232 0.069 0.013 11.275 4.127 0.072 0.027
0.875 5.600 0.087 0.019 0.125 1.878 0.028 0.036
Ablation studies. Tables 4 and 5 present some ablation studies of FAHNES. More precisely, we analyze the importance of the budget and minibatch OT-coupling. In general terms, we observe that using budgets instead of concatenating the target size to each node embedding has an important effect on the Node Num, Node Deg, and Edge Size metrics. Similarly, the minibatch OT-coupling has a positive effect on the validity of the spectral similarity of the generated hypergraphs.
We also observe additional interesting properties tied to the budget: i) training is more stable, with smoother curves for the metrics on the validation and test sets, and ii) we observe less variance in the quality of generated samples, which translates into a higher correlation between the metrics of the validation set and those of the test set. We hypothesize that the constraints introduced by the node budgets maintain the predictions in the high probability regions.
Limitations. While our node budget mechanism helps mitigate the issue of missing nodes, it does not fully solve it. On molecular generation tasks, FAHNES is less competitive due to the maturity of one-shot models in this domain: in hierarchical settings, structural errors accumulate over sequential predictions. In addition, FAHNES assumes continuous features, making it less effective for categorical data like atom or bond types—scenarios where discrete diffusion is more appropriate. Similarly, specialized techniques such as marginal probabilities [7], designed specifically for molecule generation, cannot be readily integrated into our setting. Furthermore, our framework prioritizes scalability, which limits the use of high-capacity backbones such as graph transformers [40]. Future work could explore discrete diffusion for topology and adopt a mixed discrete-continuous formulation [22] to better handle structured, symbolic domains.
@@ 5 Conclusion
We presented FAHNES, the first hierarchical framework for jointly generating hypergraph topology and features. By combining coarse-to-fine structural modeling with feature-aware generation, our method enables scalable synthesis of complex hypergraphs, surpassing the limitations of flat or disjoint approaches. Key innovations include a recursive node budget mechanism and the use of minibatch OT-coupling, which together improve sample quality and training convergence. Experiments across synthetic, molecular, and geometric datasets demonstrate strong performance in both topology and feature generation, often outperforming existing baselines. While FAHNES is not yet competitive with domain-specific pipelines—particularly for molecule generation—it opens promising directions for structured generative modeling.
Appendix A Implementation Details
A.1 Model Architecture
The denoising network is structured as follows:
  1. 1.
  2. 2.
  3. 3.
    Feature concatenation:
  4. 4.
  5. 5.
    • Edge head: Predicts edge existence.
A.2 Flow-matching Framework
$\mathcal{L}=\mathbb{E}_{t\sim\mathcal{U}(0,1),\>\mathbf{x}_{0}\sim p_{0},\>\mathbf{x}_{1}\sim p_{1}}\left[\left\|\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)-\mathbf{x}_{1}\right\|^{2}\right],$ (13)
where $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}$ is a linear interpolation between samples from the prior and target distributions.
Modeling setup. We use different prior distributions depending on the task:
  • Node and edge predictions: Prior samples $p_{0}$ are drawn from a Gaussian distribution. Targets are either $-1$ or $1$, indicating binary decisions (e.g., whether a node is expanded or an edge is retained).
  • Hyperedge expansion: Similar to node and edge predictions, with targets $-1$, 00, or $1$, encoding the number of expansions (none, one, or two).
  • Budget fractions: Prior samples $p_{0}$ are drawn from a Dirichlet distribution with concentration parameter $\alpha=1.5$, linearly mapped to $[-1,1]$ via $12x-$ 1. The target is the budget fraction of each child node, linearly mapped to $[-1,1]$ via $12x-$ 1. If a cluster is not expanded, the corresponding budget becomes $1$. Parent cluster budgets are encoded using sinusoidal positional encodings [42] (dimension 32, base frequency $10^{-4}$).
  1. i)
  2. ii)
    Find the smallest index $\rho\in\{1,\dots,K\}$ such that:
    $u_{\rho}-\frac{1}{\rho}\left(\sum_{j=1}^{\rho}u_{j}-1\right)>0.$ (15)
  3. iii)
    Compute the threshold:
    $\tau=\frac{1}{\rho}\left(\sum_{j=1}^{\rho}u_{j}-1\right).$ (16)
  4. iv)
    The projection is then:
    $\mathbf{x}^{*}=\max(\mathbf{z}-\tau,0).$ (17)
A.3 Additional Details
Definition 9 (Perturbed hypergraph expansion).
Let $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$ be a bipartite graph, and let $\mathbf{v}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ and $\mathbf{v}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$ denote the left and right cluster size vectors. For a given radius $r\in\mathbb{N}$ and probability $0\leq p\leq 1$, we construct $\tilde{B}$ as in Definition 4. Additionally, for each pair of distinct nodes $\mathbf{v}_{L}{(p)}\in\tilde{\mathcal{V}_{L}}$ and $\mathbf{v}_{R}{(q)}\in\tilde{\mathcal{V}}_{R}$ that are within a distance of at most $12r+$ 1 in $B$, we independently add an edge $e_{\{p_{i},q_{j}\}}$ to $\tilde{\mathcal{E}}$ with probability $p$.
Minibatch OT-coupling. Minibatch OT-coupling [12, 13] has demonstrated effectiveness in image generation tasks by encouraging models to learn more linear and efficient generative trajectories. Rather than sampling from the prior and target distributions independently, the two distributions are sampled jointly through the following procedure. First, samples are drawn from the prior distribution (typically Gaussian), followed by sampling corresponding targets (e.g., images). Then, the prior samples are reindexed to minimize the total distance between each input and its associated output. This coupling yields faster convergence and smoother generative flows, thereby reducing the number of sampling steps required during inference.
We extend this idea to our setting by treating the children of an expanded node as a minibatch and applying OT-coupling to the prior samples associated with these nodes. Importantly, we avoid mixing prior samples between “non-equivalent nodes”, i.e., nodes for which swapping the targets results in a different target hypergraph, as doing so changes the learned distribution. To prevent this, we restrict OT-coupling to groups of nodes that only differ with respect to their starting noise, i.e., nodes that are equivalent both topologically and in terms of features.
To manage computational complexity, we adopt the reasonable assumption that such groups of equivalent nodes correspond to the children of an expanded cluster. This assumption holds in practice when the joint topology-feature distribution is sufficiently complex. As a result, our method achieves faster convergence and requires fewer sampling steps at inference time, with only a very negligible impact on training efficiency. In our framework, expanded clusters yield two child nodes, so minibatch OT-coupling reduces to evaluating two permutations per cluster—an operation that can be efficiently parallelized using tensor operations. Algorithm 1 summarizes the process.
2:Reindexed noise samples $\{\tilde{z}_{i}\}$
3:Sample noise $\{z_{i}\}$ for each node in $B$
4:for all clusters $V^{(p)}\in B$ do
5:     if $|V^{(p)}|=1$ then
6:         No reassignment needed for singleton cluster
7:     else if $|V^{(p)}|=2$ then
8:         Let $i,j$ be the indices of the two nodes in $V^{(p)}$
9:         Let $x_{i},x_{j}$ be their corresponding targets
10:         Compute normal order cost:
$C_{\text{normal}}\leftarrow\|z_{i}-x_{i}\|^{2}+\|z_{j}-x_{j}\|^{2}$
11:         Compute swapped order cost:
$C_{\text{swap}}\leftarrow\|z_{j}-x_{i}\|^{2}+\|z_{i}-x_{j}\|^{2}$
12:         if $C_{\text{swap}}<C_{\text{normal}}$ then
13:              Swap $z_{i}\leftrightarrow z_{j}$
14:         end if
15:     end if
16:end for
17:return $\{\tilde{z}_{i}\}$ (reassigned noise samples)
Appendix B Proofs
B.1 Averaging Node Features for Clusters’ Features

### v2

@@ 4 Experiments and Results
In this section, we detail our experimental setup, covering datasets, evaluation metrics, implementation choices, baselines, results, and ablation studies. Full experimental details are provided in Appendix D. Visualizations of the generated samples are provided in Appendix I.
4.1 Datasets and Experimental Setup
Datasets. We evaluate our method on five featureless datasets: Stochastic Block Model (SBM) [24], Ego [25], Tree [26], ModelNet40 bookshelf, and ModelNet40 piano [27]. We also evaluate FAHNES on two featured hypergraphs datasets, comprising two sets of 3D meshes: Manifold40 bench and Manifold40 airplane [28]. Additionally, we adapt our method to graphs using the methodology outlined in Section 3.6 and evaluate it on two point cloud datasets sampled on the meshes of the two previous mesh datasets. For 3D mesh and point cloud datasets, node features are 3D positions, and edges or hyperedges do not have features.
@@ 4.1 Datasets and Experimental Setup
Method SBM Hypergraphs Ego Hypergraphs Tree Hypergraphs
Valid SBM $\uparrow$ Node Num $\downarrow$ Spectral $\downarrow$ Valid Ego $\uparrow$ Node Num $\downarrow$ Spectral $\downarrow$ Valid Tree $\uparrow$ Node Num $\downarrow$ Spectral $\downarrow$
HyperPA $2.5$ $\underline{0.075}$ $0.273$ $0.0$ $35.830$ $0.237$ $0.0$ $2.350$ $0.159$
VAE $0.0$ $0.375$ $0.024$ $0.0$ $47.580$ $0.133$ $0.0$ $9.700$ $0.124$
GAN $0.0$ $1.200$ $0.059$ $0.0$ $60.350$ $0.230$ $0.0$ $6.000$ $0.089$
Diffusion $0.0$ $0.150$ $0.031$ $0.0$ $\underline{4.475}$ $0.190$ $0.0$ $2.225$ $0.127$
HYGENE $\underline{65.0}$ $0.525$ $\underline{0.010}$ $\underline{90.0}$ $12.550$ $\mathbf{0.004}$ $\underline{77.5}$ $\mathbf{0.000}$ $\underline{0.012}$
FAHNES $\mathbf{87.8\vbox{\hbox{$\bm{\pm 3.1}$}}}$ $\mathbf{0.029\vbox{\hbox{$\bm{\pm 0.009}$}}}$ $\mathbf{0.006\vbox{\hbox{$\bm{\pm 0.004}$}}}$ $\mathbf{99.5\vbox{\hbox{$\bm{\pm 1.1}$}}}$ $\mathbf{0.128\vbox{\hbox{$\bm{\pm 0.171}$}}}$ $\mathbf{0.004\vbox{\hbox{$\bm{\pm 0.003}$}}}$ $\mathbf{89.7\vbox{\hbox{$\bm{\pm 6.0}$}}}$ $\mathbf{0.000\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $\mathbf{0.003\vbox{\hbox{$\bm{\pm 0.002}$}}}$
Table 2: Evaluation on ModelNet40.
Method ModelNet40 Bookshelf ModelNet40 Piano
Node Num $\downarrow$ Spectral $\downarrow$ Node Num $\downarrow$ Spectral $\downarrow$
HyperPA $8.025$ $\underline{0.048}$ $0.825$ $\underline{0.067}$
VAE $47.450$ $0.190$ $75.350$ $0.396$
GAN $\mathbf{0.000}$ $0.476$ $\mathbf{0.000}$ $0.697$
Diffusion $\mathbf{0.000}$ $0.079$ $\underline{0.050}$ $0.113$
HYGENE $69.730$ $0.068$ $42.520$ $0.117$
@@
Table 3: Evaluation on the graph point cloud datasets.
Method Airplane Point Clouds Bench Point Clouds
Chamfer Dist $\downarrow$ Spectral $\downarrow$ Ratio $\downarrow$ Chamfer Dist $\downarrow$ Spectral $\downarrow$ Ratio $\downarrow$
DiGress OOM OOM OOM OOM OOM OOM
DeFoG OOM OOM OOM OOM OOM OOM
FAHNES $\mathbf{0.094\vbox{\hbox{$\bm{\pm 0.006}$}}}$ $\mathbf{0.005\vbox{\hbox{$\bm{\pm 0.004}$}}}$ $\mathbf{67.311\vbox{\hbox{$\bm{\pm 44.903}$}}}$ $\mathbf{0.130\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $\mathbf{0.004\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $\mathbf{84.423\vbox{\hbox{$\bm{\pm 0.000}$}}}$
Metrics. For featureless hypergraphs, we follow the evaluation criteria in [10]. These include: i) structural comparison metrics such as Node Num (difference in node counts); ii) topological analysis with Spectral (maximum mean discrepancy between the spectral distributions). In scenarios where datasets enforce structural constraints, we report Valid—the percentage of generated samples satisfying those constraints. For all metrics except Valid, lower values indicate improved performance, while higher values are preferable for Valid. In the case of 3D meshes, we use ChamDist (the nearest training sample Chamfer distance), which computes the Chamfer distance between point clouds sampled from a generated sample and all training samples and outputs the minimum distance. For graph point cloud datasets, we report the nearest training sample Chamfer distance, the maximum mean discrepancy (MMD) between the spectral distributions of the generated samples and training set, and the ratio of various metrics computed for the generated samples and test set (see Appendix D for details). Detailed numerical results with additional metrics can be found in Appendix H.
Implementation details. Our implementation uses a custom flow-matching framework [22] together with a local PPGN architecture [9]. We also leverage our budget component using graph inpainting techniques, enabling the model to learn only the necessary parts of the distribution: i) budget splits are fixed to 1 for unexpanded clusters or those with a budget of $1$, ii) equal splits are enforced for expanded clusters with a budget of $2$, iii) clusters of size one are not allowed to expand, and iv) the features of non-expanded clusters are copied unchanged. Extensive implementation details are provided in Appendix E. The training and sampling procedures are explained in Appendix F. We also analyze the computational complexity in Appendix G.
Baselines. For the featureless hypergraphs, we compare FAHNES against HYGENE [10], HyperPA [29], a Variational Autoencoder (VAE) [30], a Generative Adversarial Network (GAN) [31], and a standard 2D diffusion model [32] trained on incidence matrix images, where hyperedge membership is represented by white pixels and absence by black pixels. For 3D meshes, we compare FAHNES with a sequential disjoint generation baseline, in which the hypergraph topology is generated first, followed by the feature generation. For the graph point cloud datasets, we compare FAHNES against DeFoG [33] and DiGress [6], two state-of-the-art flat models for graph generation.
4.2 Results and Discussion
Comparison with the baselines. Tables 1 and 3 show the comparisons for the featureless hypergraphs. We see that our node-budget and OT-coupling components improve the generation quality. Regarding the Manifold40 featured hypergraph dataset [^3] The full set of numerical results for the Manifold40 datasets is provided in Appendix H., FAHNES obtains a Chamfer distance of 0.073 and 0.049 for bench and airplane, respectively, while the sequential baseline reaches 0.143 and 0.117, respectively. This shows the better modeling capability of FAHNES in jointly generating topology and features in hypergraphs instead of a simple two-step approach. About the graph point cloud datasets, Table 3 shows that our hierarchical method is the only one able to perform at this scale, which shows its superior scalability. Detailed numerical results with more metrics are shown in Appendix H.
Budg.
Coup.
SBM Tree Ego
Valid $\uparrow$
NodeNum $\downarrow$
Spectral $\downarrow$
Valid $\uparrow$
NodeNum $\downarrow$
Spectral $\downarrow$
Valid $\uparrow$
NodeNum $\downarrow$
Spectral $\downarrow$
@@ 4.2 Results and Discussion
$\underline{86.7\vbox{\hbox{$\pm$6.7}}}$ $\underline{0.039\vbox{\hbox{$\pm$0.014}}}$ $\mathbf{0.006\vbox{\hbox{$\bm{\pm 0.004}$}}}$ $89.3\vbox{\hbox{$\pm$3.6}}$ $\mathbf{0.000\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $0.005\vbox{\hbox{$\pm$0.011}}$ $\underline{99.9\vbox{\hbox{$\pm$0.4}}}$ $\mathbf{0.073\vbox{\hbox{$\bm{\pm 0.050}$}}}$ $0.007\vbox{\hbox{$\pm$0.012}}$
$84.6\vbox{\hbox{$\pm$4.6}}$ $049\vbox{\hbox{$\pm$0.045}} 0.0$$007\vbox{\hbox{$\pm$0.007}}$ $\mathbf{95.6\vbox{\hbox{$\bm{\pm 3.7}$}}}$ $\mathbf{0.000\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $0.005\vbox{\hbox{$\pm$0.003}}$ $99.5\vbox{\hbox{$\pm$1.1}}$ $\underline{0.117\vbox{\hbox{$\pm$0.155}}}$ $\mathbf{0.004\vbox{\hbox{$\bm{\pm 0.003}$}}}$
@@
Table 5: Ablation studies on the node budget (Budg.) and minibatch OT-coupling (Coup.) for ModelNet and ManifoldNet datasets.
Budg.
Coup.
ModelNet Bookshelf ModelNet Piano ManifoldNet Airplane ManifoldNet Bench
NodeNum $\downarrow$
Spectral $\downarrow$
Node Num $\downarrow$
Spectral $\downarrow$ ChamDist $\downarrow$
Node Num $\downarrow$
Spectral $\downarrow$ ChamDist $\downarrow$
NodeNum $\downarrow$
Spectral $\downarrow$
[…]
$940\vbox{\hbox{$\pm$0.917}} 0.0$$032\vbox{\hbox{$\pm$0.013}}$ $3.622\vbox{\hbox{$\pm$1.822}}$ $055\vbox{\hbox{$\pm$0.040}} 0.0$$079\vbox{\hbox{$\pm$0.019}}$ $426\vbox{\hbox{$\pm$0.820}} 0.0$$14\vbox{\hbox{$\pm$0.007}} 0.0$$090\vbox{\hbox{$\pm$0.003}}$ $240\vbox{\hbox{$\pm$0.265}} 0.0$$018\vbox{\hbox{$\pm$0.009}}$
$\underline{0.265\vbox{\hbox{$\pm$0.496}}}$ $\mathbf{0.014\vbox{\hbox{$\bm{\pm 0.007}$}}}$ $\underline{3.155\vbox{\hbox{$\pm$3.637}}}$ $\mathbf{0.030\vbox{\hbox{$\bm{\pm 0.048}$}}}$ $\underline{0.050\vbox{\hbox{$\pm$0.005}}}$ $\underline{0.052\vbox{\hbox{$\pm$0.065}}}$ $\underline{0.012\vbox{\hbox{$\pm$0.005}}}$ $\underline{0.085\vbox{\hbox{$\pm$0.056}}}$ $\mathbf{0.020\vbox{\hbox{$\bm{\pm 0.032}$}}}$ $\mathbf{0.013\vbox{\hbox{$\bm{\pm 0.003}$}}}$
[…]
4.3 Ablation studies
Tables 4 and 5 present ablation studies of FAHNES for the node-budget and OT-coupling components. We observe that using budgets instead of concatenating the target size to each node embedding, like in [9, 10], improves generation quality. OT-coupling has a more nuanced effect, clearly improving quality on some datasets, like SBM or ManifoldNet Airplane, while not changing much for others. Those two components are also essential for feature generation, as lacking one of them results in much worse results, as seen by the large increase in ChamDist when one component is ablated.
4.4 Limitations
While our node budget mechanism helps mitigate the issue of missing nodes, it does not fully resolve it, and our method still struggles when generating very large hypergraphs, such as ModelNet Bookshelf and Piano, or very large graphs, such as point cloud datasets. These cases highlight that scalability remains challenging when both the number of nodes and the structural complexity grow substantially. Moreover, our framework currently assumes relatively simple feature distributions (e.g., 3D coordinates), and extending it to domains with richer or heterogeneous node and hyperedge attributes (such as categorical or multi-modal features) may require additional modeling components. Finally, although FAHNES improves stability through minibatch OT-coupling, training remains computationally expensive for large-scale datasets, which may limit applicability in resource-constrained settings.
@@ 5 Conclusion
We presented FAHNES, the first scalable hierarchical framework for jointly generating graph and hypergraph topology together with features. By integrating coarse-to-fine structural modeling with feature-aware generation, FAHNES overcomes the limitations of flat or disjoint approaches and enables the generation of complex structures at larger scales. Key innovations include a node budget mechanism, which provides fine-grained control over local growth, and minibatch OT-coupling, which improves alignment and stability during training. Extensive experiments on synthetic, 3D mesh, and point cloud datasets show that FAHNES achieves strong performance in both topology and feature generation while scaling to graph and hypergraph sizes that are out of reach for prior models. Our framework opens new directions for generative modeling of structured data, including applications in 3D geometry and circuit modeling, as well as extensions to richer feature modalities.
Appendix A Proofs
A.1 Averaging Node Features for Clusters’ Features